# A Survey of Deep Graph Learning under Distribution Shifts: from Graph Out-of-Distribution Generalization to Adaptation

### v1

@@ 5.1 Model-Centric Approaches
Fine-tuning is a widely used approach to adapt a pre-trained model during the test time. However, effectively leveraging information from the pre-trained model presents challenges, which are generally addressed by the following two techniques:
@@
Semi-supervised Fine-tuning. In this technique, the model is usually pre-trained in an unsupervised way to encode more transferable and generalizable structural information. Subsequently, during fine-tuning, task-related and domain-specific information is incorporated. A notable challenge here is the scarcity of labels in the target graphs, which can lead to overfitting. To avoid negative transfer, both G-Adapter [145] and AdapterGNN [146] employ the parameter-efficient fine-tuning (PEFT) strategy that takes bottleneck structure to alleviate overfitting by reducing the size of the tunable parameter. Alternatively, GraphControl [19] considers adaptively integrating the target information, incorporating two components: a frozen pre-trained model that takes the adjacency matrix as input and a trainable copy that takes node feature-based kernel matrix as input. The two components are connected through zero MLPs with expanding parameters, aiming to mitigate noise in target node features while gradually integrating downstream information into the pre-trained model. PROGRAM [147] builds robust connections between test data and prototypes with two key components: a prototype graph model that generates reliable pseudo-labels by leveraging relationships between prototypes and test data, and a robust self-training module that refines these pseudo-labels iteratively through consistency regularization.
@@
Self-supervised Adaptation. In this technique, task-related information is encoded into the pre-trained model, and subsequently, an unsupervised fine-tuning task on the target graph is followed. However, a major challenge arises during this unsupervised task: the model may lose its discriminative ability for the primary task or potentially learn irrelevant information. To address this issue, various techniques have been proposed to retain the model’s discriminative power. SOGA [148] and GAPGC [149] employ information-based design strategies while GT3 [14] tackles the challenge by intentional architecture design and additional regularization constraints. In SOGA, Mao et al. [148] utilized a loss that maximizes the mutual information between the inputs and outputs of the model to enhance the discriminatory power, whereas GAPGC [149] employs an adversarial pseudo-group contrast strategy to tackle over-confidence bias and mitigate the risk of capturing redundant information, which has a lower bound guarantee of the information relevant to the main task from the information bottleneck perspective. On the other hand, GT3 [14] structures the model to include two branches that share initial layers: a main task (classification) branch and a self-supervised branch, and only the self-supervised branch is tuned during test time so that the discriminative ability in the main-task branch can be maintained. Additionally, it also integrates additional regularization constraints between training and test output embeddings, hence enforcing their statistical similarity and avoiding substantial fluctuations.
@@
Other relevant studies, such as GTOT-Tuning [153], which transfers knowledge across tasks during test time, and GraphGLOW [154], which incorporates sharable components directly into the model, also provide valuable insights into addressing distribution shifts during test-time adaptation.
5.2 Data-Centric Approaches
@@
Data-centric approaches in test-time graph OOD adaptation focus on modifying the input data to effectively address distribution shifts between the training and target domains. These methodologies emphasize the reconstruction and augmentation of graph features to enhance the model’s performance without altering the pre-trained architecture.
@@
Graph Data Augmentation. Apart from reconstructing features, Jin et al. [18] introduce a graph transformation framework called GTRANS to address distribution shifts during test time. The graph transformation is modeled as injecting perturbations on the graph structure and node features, and is subsequently optimized via a parameter-free surrogate loss. Theoretical analyses are also provided to guide the selection of appropriate surrogate loss functions. Upon GTRANS, GraphCTA [15] further considers collaboratively integrating model adaptation and graph adaptation. The collaborative loop is formed by adapting the model using predictions from the node’s neighborhood, enhancing the graph through neighborhood contrastive learning, and then reintroducing the adapted graph into the model adaptation process. Additionally, the SGOOD framework [152] leverages task-agnostic substructures to enhance graph-level OOD detection by encoding their relationships into effective representations It is worth highlighting that such data-centric test-time graph OOD adaptation methods that prioritize adjusting the test data over modifying the pre-trained model can be especially beneficial when handling large-scale pre-trained models.
6 Dataset
TABLE III: A summary of graph OOD datasets. ”Avg. Nodes” and ”Avg. Edges” represent the average number of nodes and edges in each graph, respectively. ”Classes” indicates the number of target categories for classification tasks. ”Graphs” specifies the total number of graphs in each dataset. ”Scenario” refers to the primary context of generalization (G.) or adaptation (A.), while ”Shift Types” denote the types of distribution shifts, with covariate (Cov.) and concept (Con.) shifts. ”Domain Separation” identifies the factors that separate the domains. ”Field” describes the nature of the data, and ”Task” highlights the level at which the dataset can be used.
@@ 6 Dataset
DrugOOD [157]. DrugOOD focuses on OOD challenges in AI-aided drug discovery, particularly affinity prediction tasks with noisy annotations. This benchmark includes molecular graph datasets and evaluates model performance across various settings, such as Assay, Scaffold, and Size splits, where domain-specific knowledge is crucial for generalization under distribution shifts. DrugOOD primarily considers covariate shifts, thus providing insights into domain generalization for molecular graphs.
Notably, most existing graph learning benchmarks include only one type of shift. In addition to the benchmarks mentioned above, other widely-used datasets in graph learning tasks under distribution shifts include citation networks like Cora, Citeseer, and Pubmed, which are commonly used for node classification tasks. Table III summarizes the commonly used datasets for graph learning under distribution shifts, detailing their properties such as the average number of nodes, edges, classes, and graphs, as well as the applicable task types and distribution shift scenarios. These datasets cover the diversity of challenges faced in graph learning and highlights the importance of evaluating models across multiple types of distribution shifts.
7 Future Directions
@@
Distribution Shifts on Complex Graphs. In contrast to the substantial efforts dedicated to addressing distribution shifts on regular graphs, studies on more complex graph types, such as spatial, temporal, spatial-temporal, heterogeneous graphs, have received comparatively less attention [178]. Such complex graphs often exhibit diverse and dynamic patterns or involve entities and relationships of various types, introducing more intricate and nuanced distribution shifts. The comprehensive exploration and efficient mitigation of distribution shifts on complex graph types are pivotal for enhancing the capabilities of graph machine learning in broader scenarios, such as recommendation systems, healthcare systems, and traffic forecasting.
@@
8 Conclusion
In this survey, we comprehensively analyze existing methods related to deep graph learning under distribution shifts, covering three critical scenarios: graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. First, we establish problem definitions and explore different graph distribution shift types that can occur. Following this, we discuss related topicsand explain our categorization. Based on the proposed taxonomy, we systematically review techniques designed to enhance model generalizability or mitigate distribution shifts, exploring both model-based and data-centric approaches. Furthermore, we provide a summary of the datasets commonly utilized in this research area, highlighting their significance and relevance to the challenges associated with distribution shifts. Finally, we emphasize several challenges that remain and outline promising future research directions. We hope that this survey will serve as a valuable resource for researchers seeking to deepen their understanding of the current progress and trends in graph OOD learning.

### v2

@@ V-A Model-Centric Approaches
Fine-tuning is a prevalent method for adapting a pre-trained model during test time, as it allows the model to leverage knowledge learned from large-scale datasets. However, effectively utilizing information from the pre-trained model presents challenges, which are typically addressed through the following techniques.
@@
Semi-supervised Fine-tuning. This approach generally involves pre-training the model in an unsupervised manner to capture more transferable structural information. During fine-tuning, task-specific and domain-related information is incorporated to enhance adaptation. A notable challenge here is the scarcity of labels in the target graphs, which can lead to overfitting. To avoid negative transfer, both G-Adapter [143] and AdapterGNN [144] employ the parameter-efficient fine-tuning (PEFT) strategy that takes bottleneck structure to alleviate overfitting by reducing the size of the tunable parameter. Alternatively, GraphControl [19] introduces an adaptive mechanism for integrating target information. It consists of two components: a frozen pre-trained model that takes the adjacency matrix as input and a trainable copy that takes a node feature-based kernel matrix as input. These components are linked through zero MLPs with expanding parameters, designed to reduce noise in target node features while gradually incorporating downstream information into the pre-trained model. PROGRAM [145] builds robust connections between test data and prototypes with two key components: a prototype graph model that generates reliable pseudo-labels by leveraging relationships between prototypes and test data, and a robust self-training module that iteratively refines these pseudo-labels through consistency regularization.
@@
Self-supervised Adaptation. In this technique, task-related information is first encoded into the pre-trained model, followed by an unsupervised fine-tuning task on the target graph. However, a major challenge arises during this unsupervised phase: the model may lose its discriminative ability for the primary task or potentially learn irrelevant information. To address this issue, various techniques have been proposed to retain the model’s discriminative power. SOGA [146] and GAPGC [147] employ information-based design strategies while GT3 [14] tackles the challenge by intentional architecture design and additional regularization constraints. SOGA, employs a loss function that maximizes the mutual information between the inputs and outputs of the model to enhance the discriminatory power. GAPGC [147] employs an adversarial pseudo-group contrast strategy to tackle over-confidence bias and mitigate the risk of capturing redundant information, which has a lower bound guarantee of the information relevant to the main task from the information bottleneck perspective. On the other hand, GT3 [14] structures the model to include two branches that share initial layers: a main task (classification) branch and a self-supervised branchDuring test time, only the self-supervised branch is fine-tuned, ensuring that the discriminative ability of the main-task branch is preserved. Additionally, GT3 integrates additional regularization constraints between training and test-time output embeddings, hence enforcing their statistical similarity and avoiding substantial deviations.
@@
Other relevant studies, such as GTOT-Tuning [149], which focuses on transferring knowledge across tasks, and GraphGLOW [150], which incorporates sharable components directly into the model, also provide valuable insights into addressing distribution shifts during test-time adaptation.
V-B Data-Centric Approaches
@@
Data-centric approaches in test-time graph OOD adaptation focus on modifying input data rather than altering the pre-trained model. Unlike generalization techniques, which aim to improve robustness across diverse distributions, and training-time adaptation, which requires access to both source and target data, test-time adaptation operates solely on target data. These methods emphasize the reconstruction and augmentation of graph features to enhance model performance without modifying the pre-trained architecture.
@@
Graph Data Augmentation. Apart from reconstructing features, Jin et al. [18] introduce a graph transformation framework called GTRANS to address distribution shifts during test time. This framework models graph transformation as the injection of perturbations into both graph structures and node features, optimizing these transformations through a parameter-free surrogate loss. Theoretical analyses are also provided to guide the selection of appropriate surrogate loss functions. Building upon GTRANS, GraphCTA [15] further integrates model adaptation and graph adaptation in a collaborative manner. The collaborative loop is formed by adapting the model using predictions from the node’s neighborhood, enhancing the graph through neighborhood contrastive learning, and then reintroducing the adapted graph into the model adaptation process. It is worth highlighting that such data-centric test-time graph OOD adaptation methods that prioritize adjusting the test data over modifying the pre-trained model can be especially beneficial when handling large-scale pre-trained models.
VI Dataset
TABLE IV: A summary of graph OOD datasets. ”Avg. Nodes” and ”Avg. Edges” represent the average number of nodes and edges in each graph, respectively. ”Classes” indicates the number of target categories for classification tasks. ”Graphs” specifies the total number of graphs in each dataset. ”Scenario” refers to the primary context of generalization (G.) or adaptation (A.), while ”Shift Types” denote the types of distribution shifts, with covariate (Cov.) and concept (Con.) shifts. ”Domain Separation” identifies the factors that separate the domains. ”Field” describes the nature of the data, and ”Task” highlights the level at which the dataset can be used.
@@ VI Dataset
DrugOOD [153]. DrugOOD addresses OOD challenges in AI-aided drug discovery, particularly affinity prediction tasks with noisy annotations. This benchmark includes molecular graph datasets and evaluates model performance across various settings, such as Assay, Scaffold, and Size splits, where domain-specific knowledge is crucial for generalization under distribution shifts. DrugOOD primarily considers covariate shifts, thus providing insights into domain generalization for molecular graphs.
Notably, most existing graph learning benchmarks include only one type of shift. In addition to the benchmarks mentioned above, other widely-used datasets in graph learning tasks under distribution shifts include citation networks like Cora, Citeseer, and Pubmed, which are commonly used for node classification tasks. Table IV summarizes the commonly used datasets for graph learning under distribution shifts, detailing their properties such as the average number of nodes, edges, classes, and graphs, as well as the applicable task types and distribution shift scenarios. These datasets cover the diversity of challenges faced in graph learning and highlights the importance of evaluating models across multiple types of distribution shifts.
VII Future Directions
@@
Distribution Shifts on Complex Graphs. In contrast to the substantial efforts dedicated to addressing distribution shifts on regular graphs, studies on more complex graph types, such as spatial-temporal, heterogeneous graphs, have received comparatively less attention [174]. Such complex graphs often exhibit diverse and dynamic patterns or involve entities and relationships of various types, introducing more intricate and nuanced distribution shifts. The comprehensive exploration and efficient mitigation of distribution shifts on complex graph types are pivotal for enhancing the capabilities of graph machine learning in broader scenarios, such as recommendation systems, healthcare systems, and traffic forecasting.
@@
VIII Conclusion
In this survey, we comprehensively analyze existing methods related to deep graph learning under distribution shifts, covering three critical scenarios: graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. First, we establish problem definitions and explore different graph distribution shift types that can occur. Following this, we discuss related topics and explain our categorization. Based on the proposed taxonomy, we systematically review techniques aimed at enhancing model generalizability, as well as methods for training-time and test-time adaptation exploring both model-based and data-centric approaches. Furthermore, we provide a summary of the datasets commonly utilized in this research area, highlighting their significance and relevance to the challenges associated with distribution shifts. Finally, we emphasize several challenges that remain and outline promising future research directions. This survey aims to serve as a valuable resource for researchers looking to gain a deeper understanding of the progress and emerging trends in this rapidly evolving area of study.